OpenStudy (anonymous):

(sin x)(tan x cos x – cot x cos x) = 1 – 2 cos2x I did this already, but on one part my teacher is confused by it. if u want to help, tell me so i can copy paste my work!!!

OpenStudy (anonymous):

@amistre64 @campbell_st @Jaynator495 @Kainui @radar

OpenStudy (anonymous):

@KamiBug @JFraser

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

you guys wanna see my work and see if its right? My teacher was only confused at one part

OpenStudy (anonymous):

aw man u can't see the exponents cuz i used the exponent button on word, so all the EXPONENTS will be shown as blank BOXES. sorry for that!

OpenStudy (anonymous):

tan x cos x sin x – cot x cos x sin x sin⁡x/cos⁡x cos x sin x - cos⁡x/sin⁡x cos x sin x sin⁡x/cos⁡x cos x sin x - cos⁡x/sin⁡x cos x sin x Sin2 x – cos2 x Use the Identity- Cos(2x)= cos2 x – sin2 x = 2 cos2 x -1 <-part that confused her Sin2 x – cos2 x = -1(cos2 x – sin2 x) = -1(2 cos2 x -1) = -2 cos2 x +1 = 1-2 cos 2 x

OpenStudy (anonymous):

wait Ima do a screenshot! would that be better?

OpenStudy (campbell_st):

well look at the left hand side and make some simple substitutions tan = sin/cos and cot = cos/sin so you get \[\sin(x)(~~\frac{\sin(x)}{\cos(x)} \times \cos(x) - \frac{\cos(x)}{\sin(x)} \times \cos(x))\] simplify the part inside the brackets and post your solution

OpenStudy (anonymous):

so do i eliminate all the cos(x) and sin(x) that match up

OpenStudy (campbell_st):

your working makes sense

OpenStudy (anonymous):

but what about the "identity " part?

OpenStudy (anonymous):

look at the screenshot u can see it way better there

OpenStudy (campbell_st):

well I would to the bracket simplification 1st before multiplying by sin(x) it just reduces clutter

OpenStudy (anonymous):

ur talking about my work above right^^^

OpenStudy (anonymous):

can u write it out for me plzz

OpenStudy (anonymous):

idk how to simplify inside the brackets with sin x and cos x

OpenStudy (campbell_st):

so I would have \[\sin(x) ( \sin(x) - \frac{\cos^2(x)}{\sin(x)})\] now multiply and get \[\sin^2(x) - \cos^2(x)\] then I would use a substitution \[\sin^2(x) + \cos^2(x) = 1~~~~so~~~~~\sin^2(x) = 1 - \cos^2(x)\] which means you get \[(1 - \cos^2(x) - \cos^2(x)) = 1 - 2\cos^2(x)\] your method is find, just a bit long winded... and can be reduced by just using the basic identity and rewriting it... hope it makes sense.

OpenStudy (anonymous):

thanks so much!! can u help me correct 2 more plz? the next one is super short

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

#4 she said first step isnt clear

OpenStudy (anonymous):

for #5 she said yes, so....go ahead and make the substitution

OpenStudy (anonymous):

but for #5 idk what to do

OpenStudy (campbell_st):

ok... so for 4... again you need sin^2 + cos^2 = 1 cross multiply and get a common denominator and you have \[\frac{\sin(x)(1 + \cos(x) 1 - \cos(x))}{1 - \cos^2(x)}\] make the substitution in the denominator \[\frac{\sin(x)( 1 + \cos(x) 1 - \cos(x))}{\sin^2(x)}\] cancel the common denominator and simplify the numerator and its \[\frac{2}{\sin(x)} = 2 \times \frac{1}{\sin(x)} = 2 \csc(x)\]

OpenStudy (anonymous):

#5!!!! @campbell_st

OpenStudy (anonymous):

its above^

OpenStudy (campbell_st):

#5, well use the identity and replace sec^2 with 1 + tan^2 then collect like terms and see what happens.... look at the sign of the 1st term

OpenStudy (anonymous):

-1 ?

OpenStudy (anonymous):

or 1?

OpenStudy (anonymous):

it should equal 1 right, otherwise the identity wouldnt even make sense

OpenStudy (anonymous):

does this look okay Identity: sec^2 x = 1 + tan^2 x = -tan^2x + 1 + tan^2 =1

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

k thanks so muuuch

OpenStudy (anonymous):

@JFraser u came after we r already done...